# Preprint Notice

### v2

@@ Abstract
The pursuit of high-performance data transfer often focuses on raw network bandwidth, where international links of 100 Gbps or higher are frequently considered the primary enabler. While necessary, this network-centric view is incomplete, as it equates provisioned link speeds with practical, sustainable data movement capabilities across the entire edge-to-core spectrum. This paper investigates six common paradigms, ranging from network latency and TCP congestion control to host-side factors such as CPU performance and virtualization that critically impact data movement workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. We introduce the “Drainage Basin Pattern” conceptual model for reasoning about end-to-end data flow constraints across heterogeneous hardware and software components to address the fidelity gap between raw bandwidth and application-level throughput. Our findings are validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations and transcontinental production trials over 100 Gbps operational links. The results demonstrate that principal bottlenecks often reside outside the network core, and that a holistic hardware-software co-design enables consistent, predictable performance for moving data at scale and speed.
@@ I Introduction
The efficient movement of data from edge locations to core data centers and cloud resources is a foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for their raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps [1]. However, a predominant focus on raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path—including storage, host architectures, software design, and security measures—rather than by the network alone.
We identify this discrepancy as a “fidelity gap” between theoretical link capacity and actual application-level throughput. This gap is addressed by the “Drainage Basin Pattern” conceptual model (Fig. 1), which shifts the focus from isolated network optimization to systemic co-design. The efficacy of this approach is evidenced by achieving near-line-rate performance on a “high-speed” 100 Gbps transcontinental link (~84 Gbps in a pharmaceutical production environment) while enabling high-efficiency transfers at 1–10 Gbps at the edge. In this paper, we define a “fast” network as 10 Gbps up to (but excluding) 100 Gbps, and a “high-speed” network as 100 Gbps and faster.
Figure 1: The experience of moving data for most practitioners is typically limited to the source of the river, which corresponds to end-user activities, such as transferring photos and videos from mobile phones and moving spreadsheets from one folder to another. At such low data rates, operations appear simple. This may explain why efficient data transfer has historically received limited attention.
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II [2], sustained petabyte transfers [3, 4], and international deployments [5]—our work demonstrates that once a high-speed link is in place, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations [8, 14, 15] and historical transcontinental trials over 100 Gbps operational links.
@@
The core contribution of this paper is to identify and validate a set of widely held engineering beliefs—six specific paradigms—that continue to shape expectations and investments in production end-to-end data movement. This work examines why these paradigms often fall short across the full performance spectrum. We demonstrate that the path to reliable and predictable performance lies not in isolated component optimization, but in holistic, co-designed, and highly-integrated hardware and software systems.
The six paradigms we re-examine are: (1) the assumed dominance of network latency as the primary performance constraint; (2) the impact of packet loss and the perceived sufficiency of standard TCP Congestion Control Algorithms (CCA); (3) the presumed necessity of dedicated private lines for high-speed performance verification; (4) the linear expectation of a direct relationship between increased link bandwidth and application transfer rates; (5) the belief in the essential need for high-performance CPUs at every endpoint; and (6) the assumed universal utility of virtualization and cloud environments for high-throughput workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. Throughout this paper, all referenced systems are assumed to run Red Hat Enterprise Linux (RHEL) 9.6 [9] or free rebuilds such as Rocky Linux 9.6 [10].
@@ II Background / Related Work
Research programs such as the Global Research Platform (GRP) [11] have successfully promoted the orchestration of high-speed networks and dynamic path provisioning. A recent workshop [12] and similar community efforts have significantly increased the visibility and understanding of programmable 400 Gbps-class connectivity. However, a predominant focus on the network can lead to treating the data mover software and storage as auxiliary considerations. This work provides a re-balanced perspective from the angle of moving data at scale and speed.
@@ II-A Burst Buffers and Data Staging
We now introduce the burst buffer [6]—a high-speed intermediate storage layer. At the hardware level, burst buffers are built on Non-Volatile Memory express (NVMe) Solid State Drives (SSDs). In practice, this storage is formed by a dedicated software layer—such as a local file system or a scale‑out file store—that provides the underlying storage abstraction. The burst buffer serves as a staging area between production storage and the network. Originally pioneered in supercomputing to absorb data for high-performance computing (HPC) nodes at extreme velocities, we generalize this concept for moving data at scale and speed. A vendor-agnostic data movement appliance design employs Zettar zx [13] (see Table I for a summary of its features and capabilities) and burst buffers to create a predictable performance envelope. This enables sustained near-line-rate data movement between burst buffers over fast and high-speed WAN links.
Data staging, often defined simply as the movement of data between production storage and burst buffers, is a critical coordinating process. This operation must be straightforward, predictable, and highly efficient, as any delay in staging the data in or out of the buffer fundamentally negates the performance benefits of the high-speed storage component, thereby creating a new system bottleneck. Crucially, the need for this staging process arises from the fact that production storage systems are frequently constrained by low throughput, high latency, and protocol limitations, which makes them incapable of feeding high-speed WAN links directly.
To meet the need, the unified data mover zx was developed as the foundation for the holistic co-design principle central to this work. Its comprehensive and efficient architecture integrates all necessary functions— including over-distance transfer and internal data staging (between production storage and burst buffers), supporting diverse protocols (file and object), and handling both bulk and streaming transfers. This single, highly concurrent, and scale-out zx data mover manages the complete data placement workflow, from source storage through transit to destination storage. The predictable and scalable implementation of Burst Buffers and Data Staging is thus a foundational element of our Co-design Engineering Principle.
@@ II-B Bulk and Streaming Transfers
This distinction is critical because streaming transfers demand high storage throughput and low storage latency. The data mover must be of high concurrency.
@@ II-C Integrated Appliances vs Software-Centric Approaches
The network-centric approach discussed previously in Abstract and Introduction, while valuable, does not fully address the processes required for achieving lasting production value after initial deployment. Empirical evidence, including our own experience, also indicates that providing software alone—even with proven performance in validation trials [3][5], and [7]—can create significant operational challenges. Integration, tuning, and maintenance require deep, multi-disciplinary expertise, presenting a substantial total cost of ownership that many organizations’ IT departments are not equipped to bear. This is a common limitation of traditional software-centric models.
This paper demonstrates that the end-to-end data movement problem, despite its variations across workflows, can be addressed through a singular co-design engineering principle. It is this principle that forms the basis for our reexamination of the six paradigms, providing the holistic framework that reveals why these isolated beliefs are insufficient and how they can be overcome. We present a unified appliance architecture that embodies this principle across the entire data movement workflow spectrum. This co-design principle is applied across a range of hardware form-factors, from compact, low-power nodes and specialized Data Processing Units (DPUs) to certified enterprise servers from tier-1 vendors, supporting data rates from 1 Gbps up to 100 Gbps and beyond. These appliances encapsulate the necessary hardware/software co-design and tuning that was previously a manual, expert-driven process.
@@
The outcomes documented by these independent institutions [3],[4], and [7] are fully reproducible. Nevertheless, the most consistent and reliable path to reproducing these results in production is not through a software download and a custom integration project—which our experience shows can lead to significant performance variability—but through the use of pre-validated, integrated systems. This co-design principle has been implemented in commercially available appliances, with implementations ranging from regular server-based appliances from tier-1 vendors such as Hewlett Packard Enterprise (HPE) to more specialized form factors, demonstrating that reliable fast data movement is most effectively achieved through complete systems engineering [14] and [15]. The underlying principle is akin to a high-performance electric vehicle: engineered for extreme speed and efficiency, yet perfectly capable of simple, reliable operation for everyday tasks.
Despite the recognized challenges of end-to-end data movement, the prevailing industry and research approach continues to favor complex tuning methodologies and reliance on prohibitively expensive commercial testbeds or bespoke hardware. This complexity is particularly acute at the network edge (the "headwaters" of the “Drainage Basin Pattern”, Fig.1), where resource-constrained environments such as hospitals, clinics, and remote laboratories typically utilize 1–10 Gbps links. Deploying reliable, high-efficiency data transfer in these settings becomes cost-prohibitive when conventional solutions require specialized IT staff and expensive hardware.
@@ III-A The Latency Killer Paradigm
A prevalent belief in data-intensive computing is that network latency is the ultimate data movement killer. While this thinking is common among experienced IT professionals, our empirical testing and architectural analysis demonstrate that this paradigm requires re-examination. The perceived severity of latency is fundamentally governed by TCP's Bandwidth-Delay Product (BDP), which dictates the required window size for full bandwidth utilization.
@@
For years, established resources such as the U.S. DOE ESnet's fasterdata knowledge base [17] have documented methods to mitigate latency's impact on data rates. While ESnet provides valuable guidance as a network service provider, a comprehensive solution requires consideration of the entire environment, not just the network—a point we expand in Section III-C. The reproducible test results summarized in Fig. 2 and the appliance design in Fig. 3 together exemplify this holistic approach. See Table II for the Ethernet adapter’s ring buffer settings; Table III for the settings of “Tuned kernel.
@@
Section III-C provides descriptions of the testbed design. Among many uses of the testbed, it enables the testing of different latency values automatically (Fig. 1). Such a testbed was first constructed in the former Intel Swindon lab, in Swindon, United Kingdom (U.K.) Note that round-trip time (RTT), as reported by tools like ping [21], is approximately 2 × one-way latency.
@@
Figure 2: iperf3 latency sweep results obtained using two HPE DL380 Gen 11 server-based appliances (Fig. LABEL:fig:fig) on a latency simulation-capable testbed (max network speed 100 Gbps) established in the former Intel Swindon Lab, Swindon, U.K., in 2024. While default kernel networking settings (OOTB) show severe performance degradation under high latency, proper kernel tuning substantially reduces this penalty.”
@@ III-B Universal Packet Loss and TCP CCAs
This section provides a systematic examination of two interconnected paradigms to end-to-end data movement: Universal Packet Loss and CCAs. We begin by quantifying the data volumes supported by common transfer rates, then introduce a conceptual model—the “Drainage Basin Pattern”—that frames the entire spectrum of data movement workflows. Subsequently, we present empirical evidence that challenges the conventional view regarding packet loss and the criticality of TCP CCA selection in high-speed environments. A real-world case study, drawn from a Wall Street Journal report, illustrates the practical consequences of these paradigms. To contextualize the scale of these workflows, Table V quantifies the daily data volume achievable at common network speeds. Finally, we discuss why conflating different classes of workflows often leads to suboptimal outcomes.
@@
While most users experience only the “source of the river,” where consumer-grade networks may indeed have packet loss, the “high-speed backbone”—the deep main river channel—is an entirely different environment. This recognition is based on extensive experience since 2015 with 100 Gbps and faster connections, gained through collaboration with the U.S. DOE ESnet and various tier-1 organizations. Consequently, our observations confirm that in the well-engineered 100 Gbps+ Research and Education (R&E) networks discussed here—principally ESnet and Internet2 [24]—packet loss is practically nonexistent during normal operation. This observed reliability is also a key characteristic of peer R&E networks such as GÉANT [25], AARNet [26], and SingAREN [27].

### v3

@@ Abstract
The pursuit of high-performance data transfer often focuses on raw network bandwidth, where international links of 100 Gbps or higher are frequently considered the primary enabler. While necessary, this network-centric view is incomplete. It equates provisioned link speeds with practical, sustainable data movement capabilities. It is a common observation that lower than desired data rates manifest even on 10 Gbps links and commodity hardware, with higher-speed networks only amplifying their visibility. We investigate six paradigms – from network latency and TCP congestion control to host-side factors such as CPU performance and virtualization – that critically impact data movement workflows. These paradigms represent widely “accepted” engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. We introduce the “Drainage Basin Pattern” conceptual model for reasoning about end-to-end data flow constraints across heterogeneous hardware and software components at varying desired data rates to address the fidelity gap between raw bandwidth and application-level throughput. Our findings are validated through rigorous production-scale deployments, from 10 Gbps links to U.S. DOE ESnet technical evaluations and transcontinental production trials over 100 Gbps operational links. The results demonstrate that principal bottlenecks often reside outside the network core, and that a holistic hardware-software co-design enables consistent, predictable performance for moving data at scale and speed.
@@ I Introduction
The efficient movement of data from edge locations to core data centers and cloud resources is an unsolved foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps [7]. However, a predominant focus on high raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path— including storage, host architectures, software design, and security measures—rather than by the network alone. This holds true from 10 Gbps to 100 Gbps and faster.
The analyses in this paper are based on real-world, production-scale deployments, e.g., transcontinental links, edge-to-core transfers, i.e., operational environment-based. This operational grounding ensures that our insights directly inform system design and deployment practices, rather than being limited to synthetic or small-scale testbeds.
We identify this discrepancy as a “fidelity gap” between theoretical link capacity and actual application-level throughput. This gap manifests as lower than desired data rates, even on 10 Gbps links and modest commodity hardware. Higher-speed networks only amplify its effect. The “Drainage Basin Pattern” conceptual model (Fig. 1) addresses it by shifting the focus from isolated network optimization to systemic co-design. The efficacy of this approach is evidenced by achieving near-line-rate performance on a “high-speed” 100 Gbps transcontinental link (~84 Gbps in a pharmaceutical production environment). It also enables efficient 1–10 Gbps transfers at the edge. Detailed methodology and measurement results appear in Sections III and IV, where we analyze how host configurations influence end-to-end throughput across multiple deployment scenarios. In this paper, we define a “fast” network as 10 Gbps up to (but excluding) 100 Gbps, and a “high-speed” network as 100 Gbps and faster.
Figure 1: The experience of moving data for most practitioners is typically limited to the source of the river. This corresponds to end-user activities, such as transferring photos and videos from mobile phones or moving spreadsheets from one folder to another. At such low data rates, operations appear simple. This may explain why efficient data transfer has historically received limited attention, motivating the introduction of the Drainage Basin Pattern.
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II [53], sustained petabyte transfers [14, 15], and international deployments [28]—our work demonstrates that once a link is in place, whether fast or high-speed, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations [34, 33, 32], historical transcontinental trials over 100 Gbps operational links, and high-performance validations on the TaiWan Advanced Research and Education Network (TWAREN) [52]. This cross-generational validation confirms that the bottlenecks have shifted from the wide-area fabric to the internal systems architecture of the endpoints.
@@
The core contribution of this paper is the identification and validation of six widely held engineering paradigms. They continue to shape expectations, system design, and investment decisions in production end-to-end data movement. Through production-scale deployments and externally evaluated operational trials, we show why they often fail to predict achievable performance across the full edge-to-core spectrum. Reliable and predictable data movement emerges from holistic, co-designed, and tightly integrated hardware-software systems, instead of isolated optimization of individual components. The end-to-end behavior of these systems can be reasoned about using the proposed Drainage Basin Pattern.
The remainder of this paper evaluates the six paradigms through production-scale deployments and examines how the proposed principles generalize to emerging network speeds beyond 100 Gbps.
The six paradigms we re-examine are: (1) the assumed dominance of network latency as the primary performance constraint; (2) the impact of packet loss and the emphasis placed on standard TCP Congestion Control Algorithms (CCA); (3) the presumed necessity of dedicated private lines for high-speed performance verification; (4) the expectation of a direct relationship between increased link bandwidth and application transfer rates; (5) the belief in the essential need for high-performance CPUs at every endpoint; and (6) the assumed universal utility of virtualization and cloud environments for high-throughput workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. Throughout this paper, all referenced systems are assumed to run Red Hat Enterprise Linux (RHEL) 9.6 [45] or free rebuilds such as Rocky Linux 9.6 [49].
@@ II Background / Related Work
Research programs such as the Global Research Platform (GRP) [54] have successfully promoted the orchestration of high-speed networks and dynamic path provisioning. A recent workshop [55] and similar community efforts have significantly increased the visibility and understanding of programmable 400 Gbps-class connectivity. However, a predominant focus on the network can lead to treating the data mover software and storage as auxiliary considerations. This work provides a re-balanced perspective by examining data movement as a first-class systems problem [17, 21].
Rather than revisiting individual network mechanisms, we focus on how storage behavior, host architecture, and data mover design interact with high-speed networks in sustained production workflows. This systems-oriented framing shifts attention from isolated component optimization to end-to-end operational behavior.
@@ II-A Burst Buffers and Data Staging
We now introduce the burst buffer [5]—a high-speed intermediate storage layer. Originally pioneered in supercomputing to absorb data for high-performance computing (HPC) nodes at extreme velocities, we generalize this concept for sustained, wide-area data movement.
At the hardware level, burst buffers are built on Non-Volatile Memory express (NVMe) Solid State Drives (SSDs). In practice, this storage is formed by a dedicated software layer—such as a local file system or a scale‑out file store—that provides the underlying storage abstraction.
The burst buffer serves as a staging area between production storage and the network. A vendor-agnostic data movement appliance design employs a unified software data mover (Zettar zx [61]. See Table I for a summary of its features and capabilities) and burst buffers to create a predictable performance envelope that enables sustained near-line-rate data movement between burst buffers (fast ($100~\mathrm{Gbps}$) and high-speed ($\geq 100~\mathrm{Gbps}$) WAN links)
In this architecture, the burst buffer serves not only as a fast storage tier, but as a deliberate decoupling mechanism between the data mover and production storage services that are often optimized for capacity or ease of use, rather than throughput or latency. This decoupling allows each layer to operate within its effective performance regime while staging coordinates data movement across mismatched behaviors.
Data staging, often defined simply as the movement of data between production storage and burst buffers, is a critical coordinating process. This operation must be straightforward, predictable, and highly efficient, as any delay in staging fundamentally negates the performance benefits of burst buffering.
The need for staging arises because production storage systems are frequently constrained by throughput, latency, and protocol limitations, making them incapable of feeding high-speed WAN links directly.
This role becomes increasingly critical at aggregation points, where multiple modest-rate flows converge and expose the mismatch between distributed edge behavior and centralized storage services. Without effective staging, such convergence shifts the bottleneck from the network to storage, negating the benefits of high-speed connectivity.
To meet the need, the unified data mover zx was developed as the foundation for the holistic co-design principle central to this work. Its architecture integrates all necessary functions— including over-distance transfer and internal data staging (between production storage and burst buffers), supporting diverse protocols (file and object), and handling both bulk and streaming transfers. This single, concurrent, and scale-out data mover manages the complete data placement workflow, from source storage through transit to destination storage. Burst Buffers and Data Staging are thus a foundational element of our Co-design Engineering Principle.
@@ II-B Bulk and Streaming Transfers
This distinction is critical because streaming transfers demand both high storage throughput and low storage latency, stressing storage subsystems, memory hierarchies, and scheduling mechanisms in ways that differ fundamentally from bulk transfer. As a result, streaming workloads require architectures designed for sustained concurrency and flow rather than peak performance under isolated benchmarks.
Traditional High-Performance Computing (HPC) environments commonly integrate data movement into centralized batch-scheduled workflows, in which transfers are initiated, throttled, and terminated by external resource managers. While effective for coarse-grained compute allocation, this model introduces non-deterministic start times, queueing delay, and run-to-run jitter into the data path. These effects are not incidental: by treating data movement as a schedulable job rather than a continuously executing system service, centralized orchestration disrupts the temporal continuity required for sustained, near-line-rate streaming.
The architecture presented in this work eliminates this source of variability by removing centralized orchestration from the data path entirely. Data movement is executed as a decentralized, peer-to-peer service, coordinated implicitly through asynchronous buffer state rather than explicit global scheduling. In this model, transfer cadence emerges from the interaction between co-designed hardware resources, the host operating system, and local burst buffers, rather than from external queueing policies. As a result, throughput stability is bounded by physical system limits instead of orchestration-induced control latency, enabling predictable performance across heterogeneous deployment environments.
This design avoids the common mismatch between I/O demand and storage capability, which occurs when high-rate network ingress hits latency-dominated production storage services—without requiring application-level awareness or manual tuning.
@@ II-C Integrated Appliances vs Software-Centric Approaches
The network-centric approach discussed previously, while valuable, does not fully address the processes required for achieving lasting production value after initial deployment. Empirical evidence, including our own experience, also indicates that providing software alone—even with proven performance in validation trials [14, 15, 28], and [40]—can create significant operational challenges. Integration, tuning, and maintenance require deep, multi-disciplinary expertise, presenting a substantial total cost of ownership that many organizations are not equipped to sustain.
The resulting variability makes performance difficult to predict, reproduce, and operationalize across organizations and deployment environments, even when individual components perform well under controlled validation conditions.
This paper demonstrates that the end-to-end data movement problem, despite its variations across workflows, can be addressed through a singular co-design engineering principle. This work builds on extensive prior advances in high-speed networking and host tuning, and focuses on architectural principles that make such expertise deployable and repeatable across production environments. It is this principle that forms the basis for our reexamination of the six paradigms, providing the holistic framework that reveals why these isolated beliefs are insufficient and how they can be overcome. We present a unified appliance architecture that embodies this principle across the entire data movement workflow spectrum. This co-design principle is applied across a range of hardware form-factors, from compact, low-power nodes and specialized Data Processing Units (DPUs) to certified enterprise servers from tier-1 vendors, supporting data rates from 1 Gbps up to 100 Gbps and beyond. These appliances encapsulate the necessary hardware/software co-design and tuning that was previously a manual, expert-driven process.
@@
The outcomes documented by independent institutions, e.g., [14, 15, 40] are reproducible; nevertheless, our experience shows that operational reproducibility depends critically on the availability of pre-validated system configurations rather than ad-hoc assembly at each site. This co-design principle has been implemented in commercially available appliances, with implementations ranging from regular server-based units from tier-1 vendors such as Hewlett Packard Enterprise (HPE) to micro servers (Fig. 3). This demonstrates that reliable, efficient data movement is most effectively achieved through complete systems engineering [33, 32]. The underlying principle is akin to a high-performance electric vehicle (EV): engineered for extreme speed and efficiency, yet perfectly capable of simple, reliable operation for everyday tasks. Just as an EV is not merely a carriage with an electric motor swapped in, a purposely-built data movement appliance is not simply a server with powerful CPUs, NVMe SSDs, and high-bandwidth NICs; the hardware Bill of Materials (BOM) is meticulously formed to maximize operational and cost efficiency. The components are orchestrated via software to eliminate the non-deterministic bottlenecks inherent in general-purpose computing. In passing we will note that attempts to introduce legacy, non-optimized software data movers on such a deterministic platform result in performance regressions, negating the advantages of the co-design principle.
Despite the recognized challenges of end-to-end data movement, the prevailing industry and research approach continues to favor complex tuning methodologies and reliance on prohibitively expensive commercial testbeds or bespoke hardware. This complexity is particularly acute at the network edge — the "headwaters" of the “Drainage Basin Pattern”, Fig.1, where resource-constrained environments such as hospitals, clinics, and remote laboratories typically utilize 1–10 Gbps links. Deploying reliable, high-efficiency data transfer in these settings becomes cost-prohibitive when conventional solutions require specialized IT staff and expensive hardware.
@@ III-A The Latency Killer Paradigm
A prevalent belief in data-intensive computing is that network latency is the ultimate data movement killer. While this thinking is common among experienced IT professionals, our empirical testing and architectural analysis demonstrate that this paradigm requires re-examination. The perceived severity of latency is fundamentally governed by TCP's Bandwidth-Delay Product (BDP), which dictates the congestion window required for full bandwidth utilization.
@@
For years, established resources such as the U.S. DOE ESnet's fasterdata knowledge base [8] have documented methods to mitigate latency's impact on data rates. While ESnet provides valuable guidance as a network service provider, a comprehensive solution requires consideration of the entire environment, not just the network—a point we expand in Section III-C. The reproducible test results summarized in Fig. 2 and the appliance design in Fig. 3 together exemplify this holistic approach. See Table II for Ethernet adapter ring buffer settings and Table III for tuned Linux kernel parameters.
@@
Section III-C provides descriptions of the testbed design. Among many uses of the testbed, it enables the testing of different latency values automatically (Fig. 1). Such a testbed was first constructed in the former Intel Swindon lab, in Swindon, United Kingdom (U.K.). Note that round-trip time (RTT), as reported by tools like ping [36], is approximately 2 × one-way latency.
@@
Figure 2: iperf3 latency sweep results obtained using two HPE DL380 Gen 11 server-based appliances (Fig. 7) on a latency simulation-capable testbed (max network speed 100 Gbps) established in the former Intel Swindon Lab, Swindon, U.K., in 2024. While default kernel networking settings (OOTB) show severe performance degradation under high latency, proper kernel tuning substantially reduces this penalty.
@@ III-B Universal Packet Loss and TCP CCAs
This section provides a systematic examination of two interconnected paradigms to end-to-end data movement: the assumption of universal Packet Loss and TCP CCAs. We begin by quantifying the data volumes supported by common transfer rates, then introduce a conceptual model—the “Drainage Basin Pattern”—that frames the entire spectrum of data movement workflows. Subsequently, we present empirical evidence that challenges the conventional view regarding packet loss and the criticality of TCP CCA selection in high-speed environments. A real-world case study, drawn from a Wall Street Journal report, illustrates the practical consequences of these paradigms. To contextualize the scale of these workflows, Table V quantifies the daily data volume achievable at common network speeds. Finally, we discuss why conflating different classes of workflows often leads to suboptimal outcomes.
@@
While most users experience only the “source of the river,” where consumer-grade networks may indeed have packet loss, the “high-speed backbone”—the deep main river channel—is an entirely different environment. This recognition is based on extensive experience since 2015 with 100 Gbps and faster connections, gained through collaboration with the U.S. DOE ESnet and various tier-1 organizations. Consequently, our observations confirm that in the well-engineered 100 Gbps+ Research and Education (R&E) networks discussed here—principally ESnet and Internet2 [31]—packet loss is effectively negligible during normal operation. This observed reliability is also a key characteristic of peer R&E networks such as GÉANT [24], AARNet [1], and SingAREN [50].